OpenStudy (skaaterchiick):

Which of the following sequences is arithmetic? a.) 0, 3, 6, 9, 12, ... b.) 1, 4, 9, 16, 25, ... c.) 200, 100, 50, 25, ... d.) 4, 12, 36, 108, ...

imqwerty (imqwerty):

the consecutive terms of an arithmetic sequence have common difference If a, b, c, d, e is an arithmetic sequence then b-a=c-b=d-c=e-d check the option the one which follows this rule is an arithmetic sequence